OpenStudy (anonymous):

8x+4=-20 a.x-3 b.x=-2 c.x=2 d.x=3

OpenStudy (anonymous):

wht do you think it is

OpenStudy (anonymous):

i don't know wht think it is

OpenStudy (like1234):

Let's solve your equation step-by-step. 8x+4=−20 Step 1: Subtract 4 from both sides. 8x+4−4=−20−4 8x=−24 Step 2: Divide both sides by 8. 8x 8 = −24 8 x=−3 Answer: x=−3 http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

@like1234 is right

OpenStudy (anonymous):

thank u do u mind helping me with a few more